OpenStudy (anonymous):

I really need some help! How do I...solve this equation:

OpenStudy (anonymous):

\[\frac{ x-21 }{ 7 }+\frac{ x+4 }{ 4 }=x-4\]

OpenStudy (anonymous):

Multiply both sides for 28, then you have 4(x-21)+7(x+4)=28(x-4) 11x-56=28x-112 56=17x x=56/17

OpenStudy (anonymous):

Thank you!